1. Assume that any lines that appear to be tangent are tangent. 118
Answer #1

DOC = 2360°- (118° +2°) = 242 - 2° 2000 = { [ 180° - Looc] [x - 620] in O BAD .. LOD A = 180° - (LOD) = 180° - + 31° . = 2210